# Gravity-aware Grasp Generation with Implicit Grasp Mode Selection for Underactuated Hands

### v2

@@ Abstract
Learning-based grasp detectors typically assume a precision grasp, where each finger only has one contact point, and estimate the grasp probability. In this work, we propose a data generation and learning pipeline that can leverage power grasping, which has more contact points with an enveloping configuration and is robust against both positioning error and force disturbance. To train a grasp detector to prioritize power grasping while still keeping precision grasping as the secondary choice, we propose to train the network against the magnitude of disturbance in the gravity direction a grasp can resist (gravity-rejection score) rather than the binary classification of success. We also provide an efficient data generation pipeline for a dataset with gravity-rejection score annotation. In addition to thorough ablation studies, quantitative evaluation in both simulation and real-robot clarifies the significant improvement in our approach, especially when the objects are heavy.
@@ I Introduction
Power grasp [1] is a grasping mode where the fingers and palm contact the object at multiple points. Since the object is enveloped by the fingers, the grasp is robust against both initial position error and post-grasp force disturbance. Current state-of-the-art machine-learning-based grasp detectors [2, 3, 4, 5, 6, 7, 8, 9], though, only assume precision grasp [1] where each finger has only one contact point. Even though precision grasping is easier to handle in both the data generation and learning aspects, its limited contact region makes the grasp fragile. A typical failure mode of a precision grasp by a parallel-jaw gripper is that the object rotates around the axis connecting the two contact points, during which translational displacement also occurs to result in dropping the object. The small contact region also imposes a high demand for accurate hand-eye calibration and hand positioning, which is challenging for low-cost manipulators or those under low-impedance control for safety. To this end, in this paper, we introduce a data generation and learning framework to learn power grasping in addition to the common precision grasping. Figure 2 (a) illustrates a typical precision grasp and power grasp.
@@
Underactuated hands [10] have a larger number of joints than the number of actuation. This passive compliance allows the hand to adapt to various objects with non-typical geometries without explicitly controlling each joint. Since it makes power grasping easier, this work focuses on utilizing underactuated hands, or more specifically, Robotiq 2F-85 [11]. However, such mechanically automatic property makes learning power grasps challenging because whether the grasp is a power grasp or a precision grasp cannot be explicitly controlled but only decided through the interaction between the object and the hand’s link mechanism. In addition, since a power grasp requires the fingers to wrap around the object, it is not always possible due to collisions with the environment, such as the tabletop or other objects (see Fig. 1 middle and right.) To this end, we propose a neural network architecture that prioritizes power grasping if applicable but implicitly switches back to precision grasping if power grasping is not available. More specifically, while prior works [2, 3, 4, 5, 6, 7, 8, 9] only estimate the success probability of a grasp, we handle two values, namely (i) gravity-rejection score that represents the magnitude of disturbance in the gravity direction a grasp can support, which implicitly encourages power grasping since our data generation pipeline labels higher score for power grasps, and (ii) grasp validness which implicitly rejects grasps where the fingers will collide with the environment or it is too big for the hand to grasp.
@@
  • A grasp data generation pipeline for a dense annotation of grasp poses and the magnitude of disturbance in the gravity direction a grasp can support (gravity-rejection score) for both power grasping and precision grasping of underactuated hands.
  • A neural network model that prioritizes power grasping if applicable but implicitly switches back to precision grasping if power grasping is not available, by learning both gravity-rejection score and grasp validness.
  • A novel visualization approach to intuitively understand the network output. A thorough simulation evaluation and ablation study based on a novel evaluation procedure in which the grasp is performed against various object weights. A real-system quantitative evaluation that shows the significant improvement of our approach.
@@ II-A Training Data for Learned Grasp Detectors
The quality of the training data plays a critical role in any machine-learning-based system. For grasp detectors, it is common to take the synthetic approach since gathering real physics data is costly. Mahler et al. [12] labeled grasps with a score based on the robust grasp wrench space analysis [13]. Fang et al. [14] and Wang et al. [7] adopted the same metrics as [12]. Fang et al. [15] further extended [7] by adding a score based on the distance from the object’s center of mass. Qin et al. [2] first sampled grasps for each object and then projected them to cluttered scenes to reduce the computational cost, with antipodal score as the metric. Zhao et al. [3] took the same approach as [2]. Cai et al. [9] also took the two-staged approach but employed Q1 metric [16] in addition to the antipodal score. In this paper, we also adopt the two-staged approach, but it differs from other works in the grasp generation of power grasps of an underactuated hand and evaluation metrics for the gravity-rejection score.
While analytical metrics are computationally efficient, it is difficult to capture all aspects of a contact, such as contact force distribution or motion after a slip. Kappler et al. [17] reported that a simulation-based approach achieves a better performance. Zhou et al. [18] and Eppner et al. [19] followed [17] in shaking the hand in a gravity-less simulation. Breyer et al. [5] evaluated grasps by actually lifting the object in a simulation and classified its success or failure. Jiang et al. [6] also followed [5]. Huang et al. [8] further shake the hand after lifting the object to select only robust grasps. Those prior works use the simulator to get a set of valid grasps. However, such binary classification suffers from the sim-to-real gap since the absolute value of success or failure highly depends on the simulator’s contact model and parameters. In this paper, we also take a simulation-based approach, but we acquire a continuous spectrum of the gravity-rejection score to train the neural network to prioritize grasps with higher safety margins, such as power grasps. In addition, we propose an approximation to allow the two-staged approach to relax the high computation cost of the simulation-based approaches. Section III details our data generation pipeline.
@@ II-B Grasp Representation in Learned Graps Detectors
Figure 2: (a) Most works assume a precision grasp (top, bottom left) due to its simplicity. A power grasp (bottom right) allows a more robust grasp thanks to the larger number of contacts. (b) Illustration of Eq. 1 which approximate the multidimensional disturbance-rejection score to the gravity direction gravity-rejection score.
There are two major approaches to representing grasp with neural networks: sample-and-evaluate approach [20, 12, 21, 22, 3, 7] and feed-forward approach [23, 2, 4, 5, 6, 9, 8]. In the first approach, the network is only responsible for classifying and scoring the point cloud that is cropped and transformed to the hand coordinate; thus, it is extendable for power grasps. The multimodality that both power grasping and precision grasping may exist, though, prohibits using sample-efficient techniques such as sampling Darboux frame from the object surface [20, 22]. In order to avoid power grasping to induce a large computation time in the sampling stage due to the larger search space than other works, we opt for the feed-forward approach.
@@
A large number of works in 3D grasp detection [22, 21, 2, 3, 4, 7, 8, 9], take point cloud as the input. Extending the combination of point cloud representation and feed-forward network to power grasps, though, is challenging. Qin et al. [2] regressed translation offset and rotation of the hand from each point on the object surface. Cai et al. [9] used the location and normal of the points on the object surface to decide five degrees of freedom of the hand and regressed the remaining in-plane rotation. Since [2, 9] interpreted the surface point as candidates that face the hand, a set of precision grasp and power grasp poses may be assigned to a single point, leading to representation ambiguity. Sundermeyer et al. [4] and Huang et al. [8], on the other hand, interpreted the surface points as candidates of contact and reported a superior performance. Contact points of a power grasp, however, significantly differ among the object size, making explicit contact representation challenging (see Fig. 2 (a).)
Morrison et al. [23] employed a 2D fully convolutional network to infer per-pixel grasp probability and hand yaw angle from a depth image. Breyer et al. [5] extended it to 3D grasp detection with TSDF (truncated signed distance function) as the input, where a 3D fully convolutional network estimated per-voxel grasp probability and rotation in SO(3). Jiang et al. [6] followed [5] and jointly learned an implicit representation. The fully convolutional architecture is suitable to be extended for power grasps since (i) a precision grasp and a power grasp are assigned to different voxels (see Fig. 2 (a) bottom, the distance from the object differs for the two grasp modes), (ii) a precision grasp and a power grasp that grasps similar location on the object tends to have a similar rotation, thus we can expect a continuous grasp rotation field. In this work, we use the same backbone as [5], but differ in that the network is trained against both the continuous gravity-rejection score and the binary classification. Section IV details our network implementation.
@@ III Synthetic Data Generation with Gravity-Rejection Score
Given a set of scenes (in this work, 5K) with multiple objects (in this work, 1-5) placed in a cluttered way (see Fig. 3 bottom), our target is to generate a dense annotation (in this work, 1.2K on average for each scene) of grasps with both power grasp and precision grasp mode, and label the gravity-rejection score, which stands for the magnitude of disturbance in the gravity direction the grasp can support. Following the previous works [2, 3, 9], we first generate per-object grasp poses. Unlike the case of precision grasp, where grasp poses can be generated by sampling antipodal points on the object’s surface, it is not trivial to generate power grasps by a purely geometric approach. In addition, for underactuated hands, the final contact state is decided through the interaction between the object and the finger mechanism. We, therefore, take a simulation-based approach where we perturb a precision grasp and simulate the contact process by closing the hand against the object in a simulator. Through the contact, the object settles in the hand, and we record the hand’s pose relative to the object.
In order to train a network to prioritize power grasp, we need a grasp metric that takes a higher value for power grasp. While analytical metrics [16] are computationally efficient, they assume each contact is independent. For underactuated hands, though, the contacts are tightly coupled through the parallel link mechanism. Following [5, 6, 8], we also take a simulation approach for the grasp evaluation. Our approach, though, differs from theirs in two aspects: (i) instead of performing the simulation at the scene level, we perform at the object level and project the disturbance-rejection score to the gravity direction of the scenes, which drastically reduces the computation cost; (ii) instead of evaluating a grasp by a success/fail binary label, we treat them as a continuous value.
@@
At the object level, for each grasp, we apply an increasing external force on the object until it leaves the hand, and record the magnitude of the force as disturbance-rejection score. By applying the force on the object’s center of mass, we can encourage the network to prioritize grasping near the center of mass, in addition to prioritizing power grasping. As the gravity direction is unknown at this stage, we perform this evaluation in multiple directions, resulting in a $n$-dimentional score where $n$ is the number of the directions. In this work, we selected $n=6$. Considering that we employed 75 objects and generated 3K grasps per object on average, we can reduce the number of simulations to 1/24 of the cases where each grasp is evaluated at the scene level by the simulation.
@@
In order to project the $n$-dimentional disturbance-rejection score to the scene’s gravity direction to acquire the one-dimensional gravity-rejection score $f_{g}$, we take the following approximation:
@@
where $\bm{n}_{g}\in\mathbb{R}^{3}$ is the unit vector heading the direction of gravity, $f_{i}$ is the $i$-th element of the disturbance-rejection score, $\bm{n}_{i}\in\mathbb{R}^{3}$ is the unit vector heading the direction of the $i$-th disturbance, and $\epsilon$ is a small positive value. Figure 2 (b) illustrates the approximation and Fig. 3 illustrates the whole data generation pipeline.
Figure 3: Schematic of our data generation pipeline. We sample precision grasps based on the antipodal score and coarse power grasps by heuristically shifting the precision grasps (top-left). Those initial grasps are refined through grasp simulation (top-right). We apply external force in multiple directions and acquire a multi-dimensional disturbance-rejection score for each grasp. Finally, we create multi-object scenes and project per-object grasp poses to the scene to acquire the training data (bottom). The disturbance-rejection score is projected to the scene’s gravity direction to acquire the gravity-rejection score.
@@ V-A Analysis of Gravity-Rejection Score and Classification
The color bar on the right plot of the figures stands for the gravity-rejection score, which is a continuous field that takes a higher value when it’s closer to the object. This distribution encourages power grasping, as shown in Fig. 2 (a). Since the gravity-rejection score head is only trained against valid grasps, its output for the out-of-domain voxels is incorrect, e.g., it takes an even higher value inside the object or still takes a medium score at too-far-from-the-object region. The markers surrounded by a black circle represent the voxel with a higher grasp validness than the threshold. The three figures show that this classification is highly effective in rejecting those too-close or too-far grasps. In addition, in Fig. 6, grasps at $x=0$, $y=\pm 0.05$ region are rejected. Indeed, grasping this region is impossible due to the limitation of the hand’s maximum openness of 85 mm; thus, training data does not contain such grasps. Similarly, in Fig. 7 grasps at $x=0.03$, $y=0.12$ are rejected. This is also reasonable since the hand’s palm will collide with the objects to grasp this concave part.

### v3

@@ Abstract
Learning-based grasp detectors typically assume a precision grasp, where each finger only has one contact point, and estimate the grasp probability. In this work, we propose a data generation and learning pipeline that can leverage power grasping, which has more contact points with an enveloping configuration and is robust against both positioning error and force disturbance. To train a grasp detector to prioritize power grasping while still keeping precision grasping as the secondary choice, we propose to train the network against the magnitude of disturbance in the gravity direction a grasp can resist (gravity-rejection score) rather than the binary classification of success. We also provide an efficient data generation pipeline for a dataset with gravity-rejection score annotation. Evaluation in both simulation and real-robot clarifies the significant improvement in our approach, especially when the objects are heavy.
@@ I Introduction
The majority of state-of-the-art machine learning-based grasp detectors [1, 2, 3, 4, 5, 6, 7, 8] assume that each rigid finger only has one contact with a small contact region. Although such a grasp is easy to handle in both data generation and learning aspects, its limited contact region makes the grasp fragile. A typical failure mode with such grasp mode by a parallel-jaw gripper is that the object rotates around the axis connecting the two contact points, during which a translational displacement also occurs, resulting in the object dropping. The small contact region also imposes a high demand for accurate hand-eye calibration and hand positioning, which is challenging for low-cost manipulators or those under low-impedance control for safety.
Napier [9] pointed out that in order to firmly grasp an object, humans utilize the palm, and referred to such grasps as a power grip. Cutkosky [10] further extended the taxonomy of [9] in the robotics context and defined power grasp as grasps distinguished by large areas of contact between the grasped object and the surfaces of the fingers and palm, while classifying grasps where the object is held with the tips of the fingers and thumb as precision grasp. Despite the fact that there are multiple definitions of power grasp, e.g., Zhang et al. [11] redefined it as a type of grasp that its mechanism can resist passively against external forces without relying on feedback control of joint torques, and Zhang et al. [12] redefined it as a grasp with zero or less than zero connectivity, they share the same idea that utilizing multiple contacts of the intermediate finger links and palm improves grasp quality. In this paper, we introduce a data generation and learning framework to learn power grasps (with the definition of [10]) in addition to the common precision grasps, as shown in Figure 2 (a).
@@
Underactuated hands [13] have a larger number of joints than actuators. This passive compliance allows the hand to adapt to various objects with non-typical geometries without explicitly controlling each joint. Since it makes power grasping easier, this work focuses on utilizing underactuated hands, or more specifically, Robotiq 2F-85 [14]. However, such mechanically automatic property makes learning power grasping even challenging because whether the grasp is a power grasp or a precision grasp cannot be explicitly controlled but only decided through the interaction between the object and the hand’s link mechanism. In addition, since a power grasp requires the fingers to wrap around the object, which is not always possible due to collisions with the environment, such as the tabletop or other objects (see Fig. 1 middle and right.) To this end, we propose a neural network architecture that prioritizes power grasping if applicable but implicitly switches back to precision grasping if power grasping is not available. More specifically, while prior works [1, 2, 3, 4, 5, 6, 7, 8] only estimate the success probability of a grasp, we handle two values, namely (i) gravity-rejection score that represents the magnitude of disturbance in the gravity direction a grasp can support, which implicitly encourages power grasping since our data generation pipeline labels higher score for power grasps, and (ii) grasp validness which implicitly rejects grasps where the fingers will collide with the environment or it is too big for the hand to grasp.
@@
  • We propose to measure the grasp quality by gravity-rejection score, which is the magnitude of disturbance in the gravity direction a grasp can support in the simulation.
  • We provide a data generation / learning pipeline that can generate power grasps while automatically switching to precision grasping when power grasping is not available due to collision.
  • We propose to evaluate a grasp detector with varying object weight in order to examine both the detection accuracy and robustness of the detected grasp.
@@ II-A Training Data for Learned Grasp Detectors
Grasp quality metrics are key in generating synthetic training data for learning grasp detectors. An often selected approach is based on an antipodal grasp [15], adopted in [1, 2, 8]. By definition, though, it only supports antipodal grasps and is thus not applicable to power grasps with more than two contact points. Ferrari and Canny [16] proposed Q1 criteria, which is the maximum radius of a $\mathbb{R}^{6}$ sphere in the contact wrench space (GWS), with its center aligned with the origin, under the constraint that the L1 norm of the contact forces is one. Frequently referred to as $\epsilon$-metrics, it is adopted in many works such as [17, 18, 6, 19, 8, 20]. Weisz and Allen [21] discussed the robustness of the $\epsilon$-metrics against the grasp pose error. As the $\epsilon$-metrics’ assumption that each contact can exert arbitrary reaction force within the friction cone limited its application, Winkelbauer et al. [19] extended $\epsilon$-metrics to consider the torque constraint of articulated fingers. Nevertheless, the constraint of the $\epsilon$-metrics on the reaction force L1 norm prevents it from being adopted for power grasps that utilize kinematic constraint forces. To provide a quality metric for power grasps, Mirzal et al. [22] discussed the stability region of a power grasp. Zhang et al. [11] proposed a virtual work-based quality measure. However, although these works assume fully actuated hands, many robot hands that support power grasping (including ours) have an underactuated mechanism, making it hard to apply those metrics directly.
While analytical metrics are computationally efficient, it is difficult to capture all aspects of a contact, such as contact force distribution or motion after a slip. Kappler et al. [23] reported that a simulation-based approach achieves a better performance. Zhou et al. [24] and Eppner et al. [25] followed [23] by shaking the hand in a gravity-less simulation. Breyer et al. [4] and Jiang et al. [5] evaluated grasps by actually lifting the object in a simulation and classified its success or failure. Huang et al. [7] further shake the hand after lifting the object to select only robust grasps. Those prior works use the simulator to get a set of valid grasps. However, such binary classification suffers from the sim-to-real gap since the absolute value of success or failure highly depends on the simulator’s contact model and parameters. In this paper, we also take a simulation-based approach, but we acquire a continuous spectrum of the gravity-rejection score to train the neural network to prioritize grasps with higher safety margins. A similar idea can be found in Fang et al. [20], where a score based on the distance from the object’s center of mass was used, but our approach considers more diverse cases such as power grasping. In addition, we propose an approximation to allow the two-staged approach proposed by Qin et al. [1] to relax the high computation cost of the simulation-based approaches. Section III details our data generation pipeline.
@@ II-B Grasp Representation in Learned Graps Detectors
Figure 2: (a) Most works assume a precision grasp (left) due to its simplicity. A power grasp (right) allows a more robust grasp thanks to the larger number of contacts. (b) Illustration of Eq. 1 which approximates the multidimensional disturbance-rejection score to the gravity direction gravity-rejection score.
@@
Grasp representation is another key to learning grasp detectors. For 2D top-down grasps, the grasps were typically represented as the hand’s x-y location and yaw rotation on the height map [17, 26]. This approach is difficult to apply to power grasping because the hand’s insertion depth cannot be heuristically derived. For grasp detection in 3D space, Ten Pas et al. [27] first sampled points on the object surface and then searched for grasp candidates in the Darboux frame, followed by a learned grasp classifier. However, their “push” operation makes it difficult to detect both power grasps and precision grasps simultaneously. Qin et al. [1] regressed translation offset and rotation of the hand from each point on the object surface. Cai et al. [8] used the location and normal of the points on the object surface to decide five degrees of freedom of the hand and regressed the remaining rotation in the plane. Since [1, 8] interpreted the surface point as candidates facing the hand, a set of precision grasp and power grasp may be assigned to a single point, leading to representation ambiguity. Zhao et al. [2] detected the middle point of two contact points to reconstruct a grasp, thus only supporting antipodal grasps. Sundermeyer et al. [3] and Huang et al. [7] interpreted surface points as candidates of contact and reported superior performance. However, contact points of a power grasp significantly differ depending on object size, making explicit contact representation challenging (see Fig. 2 (a) where the contact points differ among different grasp modes.)
Wang et al. [6] first detected the “view” and evaluated grasps represented in the “view” coordinate with discretized hand depth and in-plane rotation. This approach is applicable to power grasping because a precision grasp and a power grasp are assigned to different discretized hand depths (see Fig. 2 (a) where the distance from the object differs for the different grasp modes). Breyer et al. [4] and Jiang et al. [5] voxelized the workspace and detected per-voxel grasp probability and hand rotation. This representation is also applicable to detecting power grasps because (i) a precision grasp and a power grasp are assigned to different voxels, (ii) a precision grasp and a power grasp that grasps a similar location on the object tend to have a similar rotation so that we can expect a continuous grasp rotation field. In this work, we follow this voxel approach due to the simple architecture and match for power grasping, but train the network with the gravity-rejection score.
@@ III Synthetic Data Generation with Gravity-Rejection Score
Given a set of scenes (in this work, 5K) with multiple objects (in this work, 1-5) placed in a cluttered arrangement (see Fig. 3 bottom), our target is to generate a dense annotation (in this work, 1.2K on average for each scene) of grasps with both power grasp and precision grasp mode, and label the gravity-rejection score, which stands for the magnitude of disturbance in the gravity direction that the grasp can support.
Following the previous works [1, 2, 8], we first generate per-object grasp poses. We first generate precision grasps by sampling antipodal points on the object surface. In order to generate power grasps, we perturb precision grasps in translation and rotation and simulate the contact process by closing the hand against the object in a gravityless simulation. Through the contact, the object settles in the hand, and we then record the hand’s pose relative to the object. This simulation-based approach is beneficial because we can avoid explicitly modeling the complex mechanism of underactuated fingers, where the final hand configuration is only decided through the interaction between the object and the spring-loaded closed-link mechanism. For each scene, we project the per-object grasp poses to the scene based on the object’s pose, and remove those that collide with the table or other objects.
In order to train a network to prioritize power grasping, we need a grasp metric that takes on a higher value for power grasps. As discussed in Sec. II-A, we take a simulation-based approach rather than an analytical one [15, 16, 22, 11]. However, unlike the case of [4, 5, 7] where the grasps are classified as success or failure, we measure the maximum disturance in the gravity direction that a grasp can support in the simulation. We call this a gravity-rejection score, which has a continuous value with the unit of [N].
@@
One drawback of the simulation-based approach is the computation time. In our case, the naive approach where each grasp in each scene is evaluated by a simulation will result in roughly 6M simulations with varying gravity. While this is not intractable, thanks to modern cloud parallel computing, this paper proposes an approximation to reduce the number of simulations. Specifically, we perform the simulation in the per-object stage rather than the per-scene stage. For each grasp, in a gravity-free simulation, we apply an increasing external force to the object until it leaves the hand and then record the magnitude of that force. By applying this force to the object’s center of mass, we can encourage the network to prioritize grasping near the center of mass, in addition to prioritizing power grasping.
Figure 3: Schematic of our data generation pipeline. We first sample antipodal grasps and randomize them (top-left). Those initial grasps are refined through grasp simulation (top-right). We apply external force in multiple directions and acquire a multi-dimensional disturbance-rejection score for each grasp. Finally, we create multi-object scenes and project per-object grasp poses to the scene to acquire the training data (bottom). The disturbance-rejection score is projected to the scene’s gravity direction to acquire the gravity-rejection score.
@@
Since the object’s pose relative to the scene’s gravity direction is unknown, we perform this evaluation in multiple directions, resulting in an $n$-dimentional disturbance-rejection score where $n$ is the number of directions (in this work, 6). In order to project the $n$-dimentional disturbance-rejection score to the scene’s gravity direction to acquire the one-dimensional gravity-rejection score $f_{g}$, we take the following approximation:
@@
where $\bm{n}_{g}\in\mathbb{R}^{3}$ is the unit vector heading the direction of gravity in a scene, $f_{i}$ is the $i$-th element of the disturbance-rejection score, $\bm{n}_{i}\in\mathbb{R}^{3}$ is the unit vector heading the direction of the $i$-th disturbance projected to the scene frame, and $\epsilon$ is a small positive value. Figure 2 (b) illustrates the approximation and Fig. 3 illustrates the whole data generation pipeline.
Thanks to the reduction in the number of simulations, we can employ accurate yet time-consuming simulation setups. Specifically, we use the hydroelastic contact model [28] implemented in Drake [29], which is highly realistic even with a concave-shaped object mesh. Simulation for all grasps for a single object typically takes roughly 12-18 hours with a single AWS m5.2xlarge instance, but it is easy to parallelize since each simulation is independent. In addition, the computation time highly depends on the settings. For example, a coarse mesh speeds up the simulation in exchange for a less accurate object shape, a larger simulation step linearly improves the simulation time in exchange for a higher chance of failure, and simply switching the contact model to the point-contact results in a more than x10 faster but much more unstable simulation.
@@ V-A Analysis of Gravity-Rejection Score and Classification
The color bar on the right plot of the figures stands for the gravity-rejection score, which is a continuous field that takes on a higher value near the object. This distribution encourages power grasping, as shown in Fig. 2 (a). Since the gravity-rejection score head is only trained against valid grasps, its output for the out-of-domain voxels is incorrect, e.g., it takes on an even higher value inside the object and a medium score at regions far removed from the object. The markers surrounded by a black circle represent the voxel with a higher grasp validness than the threshold. The three figures show that this classification is highly effective in rejecting those too-close or too-far grasps. In addition, in Fig. 6, grasps at $x=0$, $y=\pm 0.05$ region are rejected. Indeed, grasping this region is impossible due to the limitation of the hand’s maximum aperture of 85 mm; thus, training data does not contain such grasps. Similarly, in Fig. 7 grasps at $x=0.03$, $y=0.12$ are rejected. This is also reasonable since the hand’s base will collide with the objects in order to grasp this concave part.